[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 15, 2017
Christopher Mello - NOAA Federal
christopher.mello at noaa.gov
Sun Jul 16 03:40:06 UTC 2017
Joe Zajic is the contact point for goes-r ingest. He turned on the new
satellite data feed on Thursday for several sites. I sent him an email this
evening requesting he turn off the new feed. I do not know the details of
all of the channels.
The AWIPS2 satellite partition is very large and will not fill up. Jack
created a bd curve in AWIPS2 which I believe he tested against the NAWIPS
version. I will come in Sunday morning to fix the euro wave problem...an
assist any way I can with Goes-r
On Sat, Jul 15, 2017 at 10:56 PM, HSU Shift Log <nhc.tsbadmin at noaa.gov>
> *HSU Shift Log for Sat Jul 15 2017 22:56:46 GMT-0400 (EDT)*
> *Shift:* R
> *Forecaster:* E Blake
> *Technical Issues Summary:*
> *ATCF issues:*
> *N-AWIPS issues:* The microwave data that we get on NAWIPS has not been
> working since 1800Z. It came back around 22Z.
> *AWIPS-2/NCP issues:*
> *Other Techincal issues:* More satellite issues. Data was out between
> 20-22Z again, with GOES 15/13 suffering more than GOES 16. We could not
> reach Monica tonight. Data is now flowing as of 2215Z, but since this is
> the second night in a row that this has happened at nearly the same time,
> something is likely afoot. We could at least see the GOES16 data in AWIPS2.
> GOES-15 is out since 2230Z and it is now 00015Z. Dave Z says the disk is
> filling up again and he does not understand why-- he has a band-aid fix to
> the problem.
> *Miscellaneous issues:* Quite a difference between the IR temps of GOES16
> vs 15 tonight for Fernanda. Lots colder for G16-- bigger difference than
> yesterday. Hard to explain. Jack printed the images for comparison.
National Hurricane Center
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts